# Scoring, Reasoning, and Selecting the Best! Ensembling Large Language Models via a Peer-Review Process

### v2

@@ 1 Introduction
  • Similarity-based selection approach (Li et al., 2024; Guha et al., 2024; Si et al., 2023), instead, is mostly fully unsupervised (Guha et al., 2024; Si et al., 2023). These methods follow a simple principle: for a given query, select the response with the highest total similarity to all other responses. While such methods pioneered unsupervised post-hoc LLM Ensemble, their design remains coarse-grained—they rely on the naive similarity-based selection strategy (Li et al., 2024; Guha et al., 2024), along with shallow similarity measure of BLEU (Li et al., 2024) and limited informational utilization (Guha et al., 2024). Thus, the true potential of the ensemble remains largely untapped.
@@
LLM-PeerReview is built upon a transparent and interpretable framework that mimics the peer-review process, where its fully unsupervised nature provides flexibility and generalizability across various tasks and datasets. Within the methodology, we employ the LLM-as-a-Judge approach to leverage collective intelligence of LLMs for refined evaluation, moving beyond coarse metrics like BLEU, and propose a novel pivotal flipped-triple scoring trick to reduce scoring bias. Further, we provide both straightforward averaging and graphical model-based weighted methods to aggregate multiple scoring signals. Extensive experiments show that the proposed LLM-PeerReview significantly outperforms recent advanced similarity-based methods (Li et al., 2024; Guha et al., 2024) and all individual LLMs.
@@ 1.1 Related Work
LLM Ensemble, can be broadly categorized into three approaches. The first category, ensemble-before-inference approach (Shnitzer et al., 2023; Srivatsa et al., 2024; Ong et al., 2024), typically necessitates custom-labeled data to pretrain a classifier that routes each query. The mandatory pretraining phase and the dependency on labeled data are the primary inconveniences of these methods. The second category, ensemble-during-inference approach, can be subdivided into token-level (Yu et al., 2024; Huang et al., 2024; Xu et al., 2024; Zou et al., 2025; Chen et al., 2025a), span-level (Liu et al., 2024b; Xu et al., 2025), and process-level (Park et al., 2024) methods, depending on the level of granularity of the information considered in the ensemble. Among them, some methods (Zou et al., 2025; Chen et al., 2025a) require post-training (Yang et al., 2026). Lastly, as mentioned before, ensemble-after-inference approach mainly includes selection-then-regeneration-based (Lu et al., 2024) and aggregation-based methods (Li et al., 2024; Guha et al., 2024). Also, there are several approaches closely related to LLM Ensemble, such as using multiple prompts for ensemble inference (He et al., 2025).
@@ 2.5 Others: (b) Theoretical Analysis on Judging Process
LLM Ensemble Random (Guha et al., 2024) 68.4 $\pm$ 0.3 81.2 $\pm$ 1.2 52.2 $\pm$ 1.1 15.2 $\pm$ 0.6 54.2
Smoothie-Global (Guha et al., 2024) 63.0 91.5 59.8 27.6 60.5
Smoothie-Local (Guha et al., 2024) 73.6 85.5 61.8 18.3 59.8
@@ 3 Assessment: Experiment Setup
We evaluate four widely-used datasets, grouped into three categories. (1) Factual Recall: TriviaQA (Joshi et al., 2017; Guha et al., 2024) evaluates the accuracy of model responses to factual questions across various domains, including history, science, and geography; (2) Arithmetic Reasoning: GSM8k (Chen et al., 2021a; Hu et al., 2024) and MATH (Hendrycks et al., 2021; Hu et al., 2024) assess basic arithmetic and more advanced mathematical reasoning, respectively, with accuracy as the evaluation metric, focusing on correct numerical answers; (3) Instruction Following: AlpacaEval (Dubois et al., 2023; Hu et al., 2024) tests models’ ability to follow various instructions. We use GPT-4o-mini to evaluate the accuracy of model responses, assessing whether the model’s response exceeds the reference answer in the dataset. For these datasets, we employ the same versions as those used in previous studies (Guha et al., 2024; Hu et al., 2024).
@@
We use these well-established 7B models for ensemble: Llama-3.1-8B-Instruct, Mistral-7B-Instruct, Qwen2-7B-Instruct, and Qwen2.5-7B-Instruct. We compare the proposed LLM-PeerReview with the two categories of baselines. (1) Single LLMs: The four 7B-scale LLMs; (2) LLM Ensemble baselines: (i) Random (Lu et al., 2024; Guha et al., 2024) is a random-selection baseline that simply returns the response from a randomly chosen LLM in the ensemble. As one of the simplest ensemble strategies, this method has previously been applied to dialogue tasks (Lu et al., 2024); (ii) Smoothie-Global (Guha et al., 2024), Smoothie-Local (Guha et al., 2024), and Agent-Forest (Li et al., 2024) are recently proposed, strong similarity-based ensemble methods, as introduced in detail in Section 1; (iii) GaC (Yu et al., 2024) is a representative token-level ensemble-during-inference approach. It constructs a unified vocabulary that merges the individual dictionaries of LLMs. During inference, sampling is performed by observing the distributions from these models across the unified vocabulary.
@@
(1) For each individual large language model, we follow the setup of Smoothie (Guha et al., 2024), where the model responds once to each query. The responses from all models are stored for integration by the LLM Ensemble methods. (2) For the two variants of the baseline Smoothie (Guha et al., 2024), we set the number of neighbors as specified in the original paper. Agent-Forest (Li et al., 2024) does not require any hyperparameter configuration. For our method, we set the model temperature to 0 during the scoring process to eliminate suboptimal results caused by randomness. (3) All experiments were performed using 6 parallel Nvidia V100 32GB GPUs. All experiments with stochastic outputs were conducted three times.
@@ A.4 More Experiment Setup
TriviaQA (Dubois et al., 2023). The TriviaQA dataset consists of 950K question-answer pairs from 662K documents on Wikipedia and the web. It is designed for complex question answering, where answers cannot always be directly obtained through span-based methods. The dataset includes both human-verified and machine-generated pairs. We utilize the dataset from the Smoothie (Guha et al., 2024), containing 1000 randomly selected samples from the original dataset.
@@ A.5 More Experimental Results
LLM Ensemble Random 68.4 $\pm$ 0.3 81.2 $\pm$ 1.2 52.2 $\pm$ 1.1 15.2 $\pm$ 0.6 54.2
Smoothie-Global 63.0 91.5 59.8 27.6 60.5
Smoothie-Local 73.6 85.5 61.8 18.3 59.8
Agent-Forest 70.5 86.8 61.0 22.1 60.1
@@
Subsequently, in Figures A.4, A.5, A.6, and A.7, we present a case study for each corresponding dataset. Each case study illustrates the core computational process and final results of our method, as well as those of the comparison methods, Agent-Forest (Li et al., 2024) and Smoothie (Guha et al., 2024), highlighting the key computational steps and outcomes. We provide a more detailed analysis of the first two datasets, as shown in the captions; the patterns observed in the remaining two datasets are consistent with those in the first two datasets.
Figure A.4: Case study on dataset TriviaQA. Analysis: For this query, the correct answer is “Leander Club”. For each response generation, we followed the approach in  Guha et al. (2024) and applied truncation techniques under the same configuration. (i) For our method, the results of our variant, LLM-PeerReview, are shown in the Figure; also, the selected Response after truth inference remains correct, with the four scalar values obtained as: [2.39, 2.39, 1.18, 1.18]; (ii) For the baseline Agent-Forest, it focuses on “which response has the highest cumulative BLEU similarity to all other responses”. Since Responses 1 and 2 are correct, while Responses 3 and 4 are incorrect, the cumulative BLEU similarity of Response 4 with all others is the highest, so Agent-Forest incorrectly selects Response 4; (iii) For the baseline Smoothie, its two variants do not consider the four responses for the current query during the calculation. Instead, they either focus on identifying the best overall model or the responses from the most similar neighboring query. As a result, both variants of Smoothie also make errors.

### v3

@@ 1 Introduction
  • Similarity-based selection approach (Li et al., 2024; Guha et al., 2024a; Si et al., 2023), instead, is mostly fully unsupervised (Guha et al., 2024a; Si et al., 2023). These methods follow a simple principle: for a given query, select the response with the highest total similarity to all other responses. While such methods pioneered unsupervised post-hoc LLM Ensemble, their design remains coarse-grained—they rely on the naive similarity-based selection strategy (Li et al., 2024; Guha et al., 2024a), along with shallow similarity measure of BLEU (Li et al., 2024) and limited informational utilization (Guha et al., 2024a). Thus, the true potential of the ensemble remains largely untapped.
@@
LLM-PeerReview is built upon a transparent and interpretable framework that mimics the peer-review process, where its fully unsupervised nature provides flexibility and generalizability across various tasks and datasets. Within the methodology, we employ the LLM-as-a-Judge approach to leverage collective intelligence of LLMs for refined evaluation, moving beyond coarse metrics like BLEU, and propose a novel pivotal flipped-triple scoring trick to reduce scoring bias. Further, we provide both straightforward averaging and graphical model-based weighted methods to aggregate multiple scoring signals. Extensive experiments show that the proposed LLM-PeerReview significantly outperforms recent advanced similarity-based methods (Li et al., 2024; Guha et al., 2024a) and all individual LLMs.
@@ 1.1 Related Work
LLM Ensemble, can be broadly categorized into three approaches. The first category, ensemble-before-inference approach (Shnitzer et al., 2023; Srivatsa et al., 2024; Ong et al., 2024), typically necessitates custom-labeled data to pretrain a classifier that routes each query. The mandatory pretraining phase and the dependency on labeled data are the primary inconveniences of these methods. The second category, ensemble-during-inference approach, can be subdivided into token-level (Yu et al., 2024; Huang et al., 2024; Xu et al., 2024; Zou et al., 2025; Chen et al., 2025a), span-level (Liu et al., 2024b; Xu et al., 2025), and process-level (Park et al., 2024) methods, depending on the level of granularity of the information considered in the ensemble. Among them, some methods (Zou et al., 2025; Chen et al., 2025a) require post-training (Yang et al., 2026). Lastly, as mentioned before, ensemble-after-inference approach mainly includes selection-then-regeneration-based (Lu et al., 2024) and aggregation-based methods (Li et al., 2024; Guha et al., 2024a). Also, there are several approaches closely related to LLM Ensemble, such as using multiple prompts for ensemble inference (He et al., 2025).
@@ 2.5 Others: (b) Theoretical Analysis on Judging Process
LLM Ensemble Random (Guha et al., 2024a) 68.4 $\pm$ 0.3 81.2 $\pm$ 1.2 52.2 $\pm$ 1.1 15.2 $\pm$ 0.6 54.2
Smoothie-Global (Guha et al., 2024a) 63.0 91.5 59.8 27.6 60.5
Smoothie-Local (Guha et al., 2024a) 73.6 85.5 61.8 18.3 59.8
@@ 3 Assessment: Experiment Setup
We evaluate four widely-used datasets, grouped into three categories. (1) Factual Recall: TriviaQA (Joshi et al., 2017; Guha et al., 2024a) evaluates the accuracy of model responses to factual questions across various domains, including history, science, and geography; (2) Arithmetic Reasoning: GSM8k (Chen et al., 2021a; Hu et al., 2024) and MATH (Hendrycks et al., 2021; Hu et al., 2024) assess basic arithmetic and more advanced mathematical reasoning, respectively, with accuracy as the evaluation metric, focusing on correct numerical answers; (3) Instruction Following: AlpacaEval (Dubois et al., 2023; Hu et al., 2024) tests models’ ability to follow various instructions. We use GPT-4o-mini to evaluate the accuracy of model responses, assessing whether the model’s response exceeds the reference answer in the dataset. For these datasets, we employ the same versions as those used in previous studies (Guha et al., 2024a; Hu et al., 2024).
@@
We use these well-established 7B models for ensemble: Llama-3.1-8B-Instruct, Mistral-7B-Instruct, Qwen2-7B-Instruct, and Qwen2.5-7B-Instruct. We compare the proposed LLM-PeerReview with the two categories of baselines. (1) Single LLMs: The four 7B-scale LLMs; (2) LLM Ensemble baselines: (i) Random (Lu et al., 2024; Guha et al., 2024a) is a random-selection baseline that simply returns the response from a randomly chosen LLM in the ensemble. As one of the simplest ensemble strategies, this method has previously been applied to dialogue tasks (Lu et al., 2024); (ii) Smoothie-Global (Guha et al., 2024a), Smoothie-Local (Guha et al., 2024a), and Agent-Forest (Li et al., 2024) are recently proposed, strong similarity-based ensemble methods, as introduced in detail in Section 1; (iii) GaC (Yu et al., 2024) is a representative token-level ensemble-during-inference approach. It constructs a unified vocabulary that merges the individual dictionaries of LLMs. During inference, sampling is performed by observing the distributions from these models across the unified vocabulary.
@@
(1) For each individual large language model, we follow the setup of Smoothie (Guha et al., 2024a), where the model responds once to each query. The responses from all models are stored for integration by the LLM Ensemble methods. (2) For the two variants of the baseline Smoothie (Guha et al., 2024a), we set the number of neighbors as specified in the original paper. Agent-Forest (Li et al., 2024) does not require any hyperparameter configuration. For our method, we set the model temperature to 0 during the scoring process to eliminate suboptimal results caused by randomness. (3) All experiments were performed using 6 parallel Nvidia V100 32GB GPUs. All experiments with stochastic outputs were conducted three times.
@@ A.4 More Experiment Setup
TriviaQA (Dubois et al., 2023). The TriviaQA dataset consists of 950K question-answer pairs from 662K documents on Wikipedia and the web. It is designed for complex question answering, where answers cannot always be directly obtained through span-based methods. The dataset includes both human-verified and machine-generated pairs. We utilize the dataset from the Smoothie (Guha et al., 2024a), containing 1000 randomly selected samples from the original dataset.
@@ A.5 More Experimental Results
LLM Ensemble Random (Guha et al., 2024b) 68.4 $\pm$ 0.3 81.2 $\pm$ 1.2 52.2 $\pm$ 1.1 15.2 $\pm$ 0.6 54.2
Smoothie-Global (Guha et al., 2024a) 63.0 91.5 59.8 27.6 60.5
Smoothie-Local (Guha et al., 2024a) 73.6 85.5 61.8 18.3 59.8
Agent-Forest (Li et al., 2024) 70.5 86.8 61.0 22.1 60.1
GaC (Yu et al., 2024) 71.5 91.8 54.0 23.6 60.2
@@
Subsequently, in Figures A.4, A.5, A.6, and A.7, we present a case study for each corresponding dataset. Each case study illustrates the core computational process and final results of our method, as well as those of the comparison methods, Agent-Forest (Li et al., 2024) and Smoothie (Guha et al., 2024a), highlighting the key computational steps and outcomes. We provide a more detailed analysis of the first two datasets, as shown in the captions; the patterns observed in the remaining two datasets are consistent with those in the first two datasets.
Figure A.4: Case study on dataset TriviaQA. Analysis: For this query, the correct answer is “Leander Club”. For each response generation, we followed the approach in  Guha et al. (2024a) and applied truncation techniques under the same configuration. (i) For our method, the results of our variant, LLM-PeerReview, are shown in the Figure; also, the selected Response after truth inference remains correct, with the four scalar values obtained as: [2.39, 2.39, 1.18, 1.18]; (ii) For the baseline Agent-Forest, it focuses on “which response has the highest cumulative BLEU similarity to all other responses”. Since Responses 1 and 2 are correct, while Responses 3 and 4 are incorrect, the cumulative BLEU similarity of Response 4 with all others is the highest, so Agent-Forest incorrectly selects Response 4; (iii) For the baseline Smoothie, its two variants do not consider the four responses for the current query during the calculation. Instead, they either focus on identifying the best overall model or the responses from the most similar neighboring query. As a result, both variants of Smoothie also make errors.